# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v1

@@ 4.1.3 Segmentation Performance (III): False Detection Ratios, Data Loss, and Downlink Efficiency
Before concluding the evaluation of the models’ segmentation performance, we conduct a final analysis of the ranking system to evaluate its effectiveness in sorting images, as explained next. We center our analysis on the Spearman’s coefficient (Table II) to determine which model precisely ranks the HS images based on the priority levels of sea, land, and cloud coverage calculated from the segmented images. According to Tables II and LABEL:Table:models_sorted_according_to_performance, the experimental results confirm that 1D-Justo-LiuNet and 1D-Justo-HuNet are the top performers, surpassing again the state-of-the-art models. As a matter of fact, Table II shows that 1D-Justo-LiuNet and 1D-Justo-HuNet are the only models capable of reaching the ideal Spearman’s coefficient of 1.00, indicating the perfect sorting of the HS images in the rankings, without any errors compared to the ground-truth ranks for sea, land, and cloud levels. This emphasizes the selection of 1D-Justo-LiuNet (or possibly 1D-Justo-HuNet) as the preferred choice for the on-board ranking system.
@@
Furthermore, the results reveal that our 2D model, 2D-Justo-UNet-Simple, although not achieving the same level of ranking performance as the 1D models, also surpass most of their fellow 2D models from the state of the art. Nonetheless, our model, 2D-Justo-LucasCNN, does not produce improved results, as it only yields the same figures to those achieved by other state-of-the-art models (1D-ML/SGD, 2D-CUNet Reduced, and 2D-UNet FAUBAI).
4.1.4 Segmentation Performance (IV): Summary
@@
In light of the previous sections, in addition to the key finding that 1D models outperform their 2D counterparts, we conclude that our 1D-Justo-LiuNet model consistently outperforms the current state-of-the-art models, all trained for the same sea-land-cloud segmentation task in HYPSO-1 data. Specifically, the 1D-Justo-LiuNet model surpasses other models in the literature in terms of segmentation performance, assessed by average accuracy and false detection ratios for each class. Furthermore, these superior results for the 1D-Justo-LiuNet are consistent even in the context of the ranking system application case, and also when compared with the soil texture literature [42]. However, we stress that our results and conclusions are specific to the sea-land-cloud dataset [31], and we acknowledge that model performance may differ when the data source is altered in any manner. For instance, a model that performs well on one dataset may not excel on another. Consequently, to facilitate testing of the models used in this work on other data sources, we provide the software code for all models in the supplementary materials accompanying this paper.
@@ 4.1.5 Parameter Count
Following the analysis of the models’ performance, we next examine the number of model parameters that need to be uplinked to a satellite for in-orbit inference (Table II). Our lightweight model 1D-Justo-LiuNet surpasses the existing literature in terms of parameter count (Table LABEL:Table:models_sorted_according_to_performance), ranking at the top, as the model with the fewest parameters, even when compared to other lightweight models in the literature, such as 2D-CUNet, 2D-CUNet++, 2D-CUNet Reduced, and 2D-CUNet++ Reduced [38]. Hence, the lightweight 1D-Justo-LiuNet model stands out as the most practical choice when solely considering the reduction in on-board memory required for storing model parameters. Moreover, despite some of our other models, such as 1D-Justo-HuNet and 2D-Justo-UNet-Simple, exhibit a slight increase in the number of parameters compared to 1D-Justo-LiuNet, they still maintain a parameter count below ten thousands, which is still lower than the parameters of the remaining state-of-the-art models. As mentioned earlier, the lowest-ranked 2D-UNet-FAUBAI network utilizes over 26 million parameters. Even its reduced version, 2D-UNet-FAUBAI-reduced, has nearly 2 million parameters, which we find excessive for on-board deployment while it does not improve the model’s performance neither in terms of segmentation accuracy nor in terms of image ranking. Finally, we emphasize that a significant majority of our top models in terms of performance (accuracy and Spearman’s coefficient in Table LABEL:Table:models_sorted_according_to_performance), occupy the best ranks in terms of parameter count. Therefore, we can assert that we achieve the objective outlined in Section 2 regarding surpassing the state-of-the-art performance, while significantly compressing the models.
@@ 4.1.6 Inference Time
The final metric of our analysis is the average inference time per HS image (Table II). Despite the strengths of the 1D models, with particular emphasis on the favorable results by the 1D-Justo-LiuNet, 1D models exhibit a drawback of extended inference times on the processing architecture used in the tests, even considering their lowest parameter count. In Table II, we observe that both 1D ML and DL models, especially DL models with increased intrinsic complexity, exhibit lengthier inference times. In fact, the results reveal that the 1D-Justo-LiuNet has the lengthiest inference time, at around 15 seconds for a single HS image. The results demonstrate, however, that 2D models significantly decrease the inference time, reducing it to the order of milliseconds. The disparity in speed may be attributed to the GPU being primarily optimized for high parallelization of 3D imaging data, which is a prevalent format in computer vision. Consequently, GPUs may exhibit greater efficiency when processing 3D image patches as opposed to 1D sequential data (spectral signatures).
@@
Although satellites like HYPSO-1 experience data latency lasting minutes to hours [26], we propose the following solution if inference times in the order of seconds (e.g., 1D-Justo-LiuNet) are perceived as lengthy for in-orbit inference. Employing the HYPSO-1 satellite as an example case, its On-board Processing Unit (OPU) integrates an FPGA System-on-Chip (Zynq-7030/XC7Z030-1) [9]. This system features the ARM Cortex-A9 processor and provides superior FPGA resources in comparison to other platforms, such as the Zynq-7020, previously employed in cloud detection within satellite imagery [13]. As future work, we suggest implementing the 1D-Justo-LiuNet in software on the processor side to confirm whether the inference time is still lengthy. We anticipate that the processor’s lower-level implementation may yield shorter inference times compared to running inference in higher-level languages like Python. If the inference time would still remain too long, we recommend a combined software-hardware co-design approach. This approach accelerates computationally intensive tasks for faster inference, while maintaining low power consumption using the FPGA fabric logic.
4.2 Analysis of Results for Models Trained on 3 Spectral Channels
@@
We finalize the discussion section by addressing the impact of reducing the number of channels from 112 to 3 on the reported metrics. As indicated in Table IV, when we retrain the models using the 3-channel dataset instead, the inference time decreases compared to the models trained with 112 channels, as there is now less data to process. Moreover, there is a reduction in the number of model parameters due to the smaller number of components within each neural kernel used for computing the corresponding 2D convolutions across the spectral dimension. Nevertheless, the reduction in feature channels leads to a decreased segmentation performance, with the highest accuracy reaching 0.87, unlike the 0.93 top accuracy for the models trained with 112 channels. This decline in performance is also evident in the evaluation of the ranking system, for which none of the models achieve the ideal coefficient of 1.00, as observed earlier with 112 channels. As a matter of fact, the highest Spearman’s coefficient is notably lower at 0.83. Although the main emphasis of our analysis is to examine the impact of reducing the channels from 112 to 3, we shortly discuss next how the 2D models compare to each other for solely 3 channels instead. Table IV shows that the 2D-Justo-Unet-Simple model achieves an accuracy of 0.85, which is close to the highest accuracy of 0.87. In the ranking system, our model exhibits the same top Spearman’s coefficient as 2D-CUNet and 2D-CUNet++ Reduced. Remarkably, the 2D-Justo-UNet-Simple outperforms both 2D-CUNet and 2D-CUNet++ Reduced in terms of both parameter count and inference time.
@@ 5 Conclusion
In this work, we conducted experiments including the training and comparison of 16 ML and DL models for on-board deployment using the Sea-Land-Cloud-Labeled dataset from the HYPSO-1 mission [31], as an illustrative case. We demonstrated that for segmenting sea, land, and cloud categories in hyperspectral satellite imagery, 1D models surpass 2D models in terms of performance. This observation also holds for our proposed ranking system combining decision-making for sea, land and cloud coverage. Our model 1D-Justo-LiuNet consistently delivers better results than the state-of-the-art models. Furthermore, we also found that using L1b calibrated radiance data, instead of raw captured data, improved segmentation performance. Moreover, training models with only three spectral channels led to shorter inference times and reduced parameter counts, at the cost of lower segmentation performance. Finally, as future work, we suggest conducting experimental tests to deploy in orbit the 1D-Justo-LiuNet model.
@@
The labeled dataset to reproduce the findings in this paper is openly available to download at https://ntnu-smallsat-lab.github.io/hypso1_sea_land_clouds_dataset/. Moreover, we provide as supplementary materials: software codes for the models, trained model files with the learned parameters, information on training-validation-testing, additional evaluation metrics, and generalization tests. The material can be found in our GitHub repository at https://github.com/jonalvjusto/s_l_c_segm_hyp_img.
@@ Acknowledgements
The research leading to these results has received funding from the NO Grants 2014 – 2021, under Project ELO-Hyp contract no. 24/2020, and the Research Council of Norway grant no. 223254 (AMOS), no. 328724 (Green Platform) and no. 325961 (HYPSCI). We extend our gratitude to Anum Masood, Per G. Kjeldsberg, and Simen Netteland.

### v2

@@ 4.1.4 Segmentation Performance (IV): Priority Rankings for Downlink
As part of the further evaluation of the models’ performance, we conduct an additional analysis of the ranking system to evaluate its effectiveness in sorting HS images, as explained next. We center our analysis on the Spearman’s coefficient (Table II) to determine which model precisely ranks the images based on the sea, land, and cloud coverage levels calculated from the segmented images. According to Tables II and LABEL:Table:models_sorted_according_to_performance, the experimental results confirm that 1D-Justo-LiuNet and 1D-Justo-HuNet are the top performers, surpassing again the state-of-the-art models. As a matter of fact, Table II shows that 1D-Justo-LiuNet and 1D-Justo-HuNet are the only models capable of reaching the ideal Spearman’s coefficient of 1.00, indicating the perfect sorting of the images in the rankings, with no errors compared to the ground-truth ranks for sea, land, and cloud levels. This emphasizes the selection of 1D-Justo-LiuNet (or possibly 1D-Justo-HuNet) as the preferred choice for the proposed on-board ranking system.
@@
Furthermore, the results reveal that our 2D model, 2D-Justo-UNet-Simple, although not achieving the same level of ranking performance as the 1D models, also surpass most of their fellow 2D models from the state of the art. Nonetheless, our model, 1D-Justo-LucasCNN, does not produce improved results, as it only yields the same figures to those achieved by other state-of-the-art models (1D-ML/SGD, 2D-CUNet Reduced, and 2D-UNet FAUBAI).
4.1.5 Segmentation Performance (V): Model Generalisation in Deployment Set
Aligning with the segmentation literature for satellite imagery [13], we conduct generalization tests in the unseen on-ground deployment set of 30 images, relying on expert visual inspection (similar to [13]) as this set has no ground-truth labels available. From the generalization tests, we next illustrate the segmentation results for one of the deployment images, specifically a HYPSO-1 capture taken over the Namib desert in Africa [34], as depicted in Fig. 6. Fig. 6(a) presents the segmentation results obtained using 1D-Justo-LiuNet. The figure includes a grayscale L1b radiance image at a wavelength of $\lambda$=699.61 nm, followed by the model’s segmented image, a confidence probability map indicating the model’s certainty for each pixel-level prediction, and summarised information regarding sea, land, and cloud coverage levels used to sort the images in the ranking system. The figure illustrates that 1D-Justo-LiuNet performs adequately, with a slightly reduced certainty in class transitions while maintaining a high level of overall certainty in its predictions. Regarding 2D-CNN models, due to the limited generalization capability observed in the deployment set for 2D-Justo-UNet-Simple (which is, however, a top-performer in the test set), we choose to conduct generalization tests on other 2D-CNN models of interest instead, such as 2D-CUNet++ Reduced, as illustrated in Fig. 6(b). The figure highlights that 2D-CUNet++ Reduced encounters important misclassifications at the non-overlapping patch borders, consistent with the confidence observed in the patch edges in addition to the decreased certainty in the class transitions. This phenomenon may be attributed to a variety of factors, including the network’s padding of input patches, among other potential causes.
In the supplementary material for the generalisation tests on the rest of the deployment set, we provide the outcomes for the remaining deployment images. From the overall on-ground deployment, we conclude that 1D-Justo-LiuNet generalizes acceptably for our needs. However, its 1D-CNN architecture can result in predictions with higher entropy, leading to less smooth class transitions. This is attributed to the absence of spatial context in its pixel-level predictions. On the other hand, 2D-CUNet++ Reduced, while still generalizing effectively, exhibits higher misclassifications at patch borders. Nevertheless, it remarkably provides predictions with lower entropy, resulting in smoother class transitions, due to the utilization of spatial neighboring information during pixel-level predictions.
4.1.6 Segmentation Performance (VI): Summary
@@
In light of the previous sections, in addition to the key finding that 1D models outperform their 2D counterparts in our experiments, we conclude that the 1D-Justo-LiuNet model consistently outperforms the current state-of-the-art models, all trained for the same sea-land-cloud segmentation task in HYPSO-1 L1b calibrated data. Specifically, the 1D-Justo-LiuNet surpasses the literature in terms of segmentation performance, assessed by average accuracy and false detection ratios trade-off. Furthermore, our results for the 1D-Justo-LiuNet not only surpass the performance reported by the soil spectroscopy literature [45], but our findings also exhibit remarkable consistency, even within the context of the ranking system application case (Spearman’s coefficient), where the 1D-Justo-LiuNet also outperforms the other models. Additionally, 1D-Justo-LiuNet demonstrates acceptable performance when subjected to generalization tests on the unseen on-ground deployment set. However, we stress that our results and conclusions are specific to the sea-land-cloud dataset [34], and we acknowledge that model performance may differ when the data source is altered in any manner. For instance, a model that performs well on one dataset may not excel on another. To facilitate testing of the models used in this work on other data sources, we provide the software code for all models in the supplementary materials accompanying this paper.
@@ 4.1.7 Parameter Count
Following the analysis of the models’ performance, we next examine the number of model parameters (Table II) that need to be uplinked to a satellite for in-orbit deployment. Our lightweight model 1D-Justo-LiuNet surpasses the existing literature in terms of parameter count ranking at the top (Table LABEL:Table:models_sorted_according_to_performance), as the model with the fewest parameters (4,563), even when compared to other lightweight models in the literature, such as 2D-CUNet, 2D-CUNet++, 2D-CUNet Reduced, and 2D-CUNet++ Reduced. Hence, the lightweight 1D-Justo-LiuNet model stands out as the most practical choice when solely considering the reduction in on-board memory required for storing the model parameters. Moreover, despite some of our other models, such as 1D-Justo-HuNet and 2D-Justo-UNet-Simple, exhibit a slight increase in the number of parameters compared to 1D-Justo-LiuNet, they still maintain their parameter count below ten thousands, which is still lower than the number of parameters of the other state-of-the-art models. As mentioned earlier, the lowest-ranked 2D-UNet FAUBAI network utilizes over 26 million parameters. Even its reduced version, 2D-UNet FAUBAI reduced, has nearly 2 million parameters, which we also find excessive for on-board deployment while it does not improve the model’s performance neither in terms of segmentation accuracy nor in terms of image ranking. Finally, we emphasize that a significant majority of our top models in terms of performance (accuracy and Spearman’s coefficient in Table LABEL:Table:models_sorted_according_to_performance), also occupy the best ranks in terms of parameter count. Therefore, we can assert that we achieve our objective outlined in Section 2 regarding surpassing the state-of-the-art performance, while significantly compressing the models.
@@ 4.1.8 Inference Time
The final metric of our analysis is the average inference time per HS image (Table II). Despite the strengths of the 1D models, with particular emphasis on the favorable results by the 1D-Justo-LiuNet, 1D models exhibit a drawback of extended inference times on the processing architecture used in the tests, even considering their low parameter count. In Table II, we observe that both 1D ML and DL models, especially DL models with more intrinsic complexity, exhibit lengthier inference times. In fact, the results reveal that the 1D-Justo-LiuNet has the lengthiest inference time, at around 15 seconds for a single HS image. The results demonstrate, however, that 2D models significantly decrease the inference time, reducing it to the order of milliseconds. The disparity in speed may be attributed due to the potentially greater number of floating-point operations per second (FLOPS) required by the 1D models, among other factors affecting the inference time.
@@
Although satellites like HYPSO-1 experience data latency lasting minutes to hours [29], we propose the following solution if inference times in the order of seconds (e.g., by 1D-Justo-LiuNet) are perceived as lengthy for in-orbit inference. Employing the HYPSO-1 satellite as an example case, its On-board Processing Unit (OPU) integrates an FPGA System-on-Chip (Zynq-7030/XC7Z030-1) [10]. The system features an ARM Cortex-A9 processor and provides superior FPGA logic resources in comparison to other platforms, such as the Zynq-7020, previously employed in cloud detection within satellite imagery [14]. As future work, we suggest implementing the 1D-Justo-LiuNet in the processor side to confirm whether the inference time is still lengthy. We anticipate that the processor’s lower-level software implementation may yield shorter inference times compared to running inference in higher-level languages like Python. If the inference time would still remain too long, we recommend a combined software-hardware co-design approach. This approach accelerates computationally intensive tasks for faster inference, while maintaining low power consumption, by using the FPGA fabric logic.
4.2 Results Analysis for Models Trained on 3 Channels
@@
We finalize the discussion section by addressing the impact of reducing the number of channels from 112 to 3 on the reported metrics. As indicated in Table IV, when we train the models using the 3-channel dataset instead, the inference times decrease compared to the models trained with 112 channels, as there is now less data to process. Moreover, there is a reduction in the number of model parameters due to the smaller number of components within each convolutional kernel when computing the respective 2D convolutions across the reduced spectral dimension. Nevertheless, the reduction in feature channels leads to a decreased segmentation performance, with the highest accuracy reaching 0.87, unlike the 0.93 top accuracy for the models trained with 112 channels. The decline in performance is also evident in the evaluation of the ranking system, for which none of the models achieve the ideal coefficient of 1.00, as observed earlier with 112 channels. As a matter of fact, the highest Spearman’s coefficient for 3 channels is notably lower at 0.83.
@@ 5 Conclusion
In this work, we conducted experiments including the training and comparison of 16 machine learning and deep learning models, whose software codes are provided within this article along with additional supplementary materials, for future in-orbit deployment using the Sea-Land-Cloud-Labeled dataset from the HYPSO-1 mission [34], as an illustrative case. Our experiments demonstrated that when segmenting sea, land, and cloud categories in hyperspectral satellite imagery, 1D models surpass 2D models in terms of performance. This conclusion also holds for our proposed ranking system combining priority downlink according to sea, land and cloud coverage. Our proposed lightweight models, such as 1D-Justo-LiuNet and 1D-Justo-HuNet, consistently surpass the performance of state-of-the-art models. Moreover, they demonstrate remarkable model compression, reaching up to 90.50% when compared to the existing models. This makes 1D-Justo-LiuNet and 1D-Justo-HuNet well-suited for future in-flight deployment. Furthermore, we also found that using L1b calibrated radiance data, instead of raw sensor data, improved segmentation performance. Moreover, training models with only three spectral channels led to shorter inference times and reduced parameter counts, at the cost of lower segmentation performance. Finally, as future work, we suggest to study the implementation of processing architectures that are used aboard satellites and conducting experimental tests to deploy, for instance, the 1D-Justo-LiuNet model in orbit.
@@
To begin, the labeled dataset used in this paper is openly available to download at https://ntnu-smallsat-lab.github.io/hypso1_sea_land_clouds_dataset/. Our supplementary material is openly available at https://github.com/NTNU-SmallSat-Lab/s_l_c_segm_hyp_img/. We provide the following resources. Firstly, the Python model codes accompanied by a notebook that illustrates how to perform inference on HYPSO-1 data, by 1D-CNN and 2D-CNN models. Secondly, there is an executable software program designed to serve as an oceanic-terrestrial-cloud segmenter for hyperspectral imagery acquired by the HYPSO-1 satellite. Thirdly, we provide trained models with learned parameters, with 46 distinct configurations, for the HYPSO-1 data. Fourth, we offer additional in-depth architectural details for the models. Fifth, we include information regarding the training, validation, and test data splits. Sixth, we share an extensive set of experimental figures, including multiple metrics. Seventh, we provide a deployment set consisting of 30 unlabeled HYPSO-1 images. Lastly, we present the generalisation tests in the complete deployment set.
@@ Acknowledgements
The research leading to these results has received funding from the NO Grants 2014 – 2021, under Project ELO-Hyp contract no. 24/2020, and the Research Council of Norway grant no. 328724 (Green Platform) and no. 325961 (HYPSCI). We extend our gratitude to Anum Masood, Per Gunnar Kjeldsberg, and Simen Netteland.
Appendix A Direct Channel Reduction
Considering that the legacy algorithms for cloud detection in HS satellite imagery typically select 3 channels to mimic RGB image processing [11, 12, 13, 14], we also obtain another dataset by reducing the 112 channels to merely 3 channels. To achieve this reduction, we explore the literature of dimensionality reduction extensively studied for HS satellite missions [46, 47, 48]. First, the focus on reducing dimensions originates from the importance of having low-dimensional data which can save computation resources, memory usage and power. Next, some of the techniques proposed by the literature include methods like Principal Component Analysis (PCA) [49] and Independent Component Analysis (ICA) [50]. In alignment with the state-of-the-art techniques, we employ PCA to reduce the 112 channels (normalized) to 3. To alleviate to a certain extent the computational demands of satellites operating often with tight power constraints, we propose to conduct both the training and inference of PCA exclusively on the ground. Eliminating PCA inference on the satellite enables potential power savings, which could be of importance to conduct the subsequent segmentation inference during flight. Consequently, and as also outlined by the literature [12], we suggest implementing on board a direct channel selection (3 channels) instead of deploying PCA inference. In this work, the purpose of employing PCA only on ground is to identify in advance the channels containing the most useful information in the data. In particular, we use PCA to identify the channels that have the greatest influence on a single principal component, as determined by the weight loadings of those channels. In the case of the 112-channel L1b dataset, we find that the majority of influential channels are predominantly within the NIR range. However, given the necessity of considering other spectral ranges in addition to NIR for the adequate segmentation of sea, land, and clouds – as the NIR range alone is likely to be insufficient – we opt to pick the top three influential channels from only the visible blue, second the visible green and red combined, and finally the NIR. In the dataset, the most influential channels within each spectral range are: 412.72 nm (channel 7) for the blue spectrum, 699.61 nm (channel 89) in the green and red spectra, and 747.77 nm (channel 103) in the NIR spectrum. These align closely with the L1b radiance bands frequently analyzed during HYPSO-1 operations, especially for channels 89 and 103.
Appendix B Data Dimensions Arrangement
On the one hand, in the case of 1D-CNN or 1D ML models, we transform each HS image into a 2D array resembling tabular data. Therefore, the models take inputs formatted as BATCH SIZE $\times$ CHANNELS, where BATCH SIZE indicates the number of image pixels for which predictions are to be inferred. 1D-CNN models perform 1D convolutions separately on each pixel across its channels. On the other hand, for 2D-CNNs, optimizing memory and computational resources during training and inference is even more essential, and to help achieving this, we divide each HS image datacube into smaller cubes, often referred as 3D patches or tiles. Padding extends the image borders to retain the integrity of edge information, and the padded borders can be eliminated by cropping the extended spatial dimensions in the segmented image. After padding, we proceed with non-overlapping patching. Unlike overlapping patching, this approach guarantees that each pixel is processed only once to reduce computational complexity. The input format for the 2D-CNNs is BATCH SIZE $\times$ PATCH SIZE $\times$ PATCH SIZE $\times$ CHANNELS, where BATCH SIZE refers in this case to the number of 3D patches. We find that a PATCH SIZE of 48 is suitable for our experiments as it aligns with the encoder-decoder networks, which repeatedly decrease the spatial dimensions by a factor of two via pooling. Consequently, selecting a PATCH SIZE of 48 ensures that the spatial dimensions subjected to pooling remain integer as the network’s depth increases while reducing memory and computational resources in forward propagation.